# thanks: ORCID 0000-0002-9656-9570thanks: ORCID 0000-0001-9896-1018thanks: ORCID 0000-0001-9980-0363thanks: ORCID 0000-0003-1168-7883

### v1

@@ Abstract
The concept of the applied field is useful in social psychology [13, 14], as it can represent incentives to improve or deteriorate interpersonal relationships. An external field in this role has been integrated into the Heider model in Refs. 15, 16. The field and thermal noise complement the toolbox of statistical mechanics [17]. Our aim here is to demonstrate the equivalence of the balanced system in the presence of an external field and the standard Ising system without a field.
@@
A balanced network has the following topological property [2]: the parity of any path $\gamma=(a_{1}a_{2},a_{2}a_{3},\ldots,a_{n}a_{n+1})$, defined as the product of the signs of the edges that form the path
$\pi(\gamma)=\prod_{i=1}^{n}a_{i}a_{i+1},$ (1)
@@
where $n_{t}=N_{t}/N=(N-1)(N-2)/6$ is the density of triangles per node, and $n_{e}=N_{e}/N=(N-1)/2$ is the density of edges per node. The first sum runs over all triangles, and the second one runs over the links of the network. In physical terms, $M$ can be called the edge magnetization, and $m=M/N$ the magnetization density. Positive/negative $M$ indicates friendly/unfriendly relations. The term $P$ sums the parities of the elementary cycles; $p=P/N$ is the average parity per cycle. There are two coupling constants $\varepsilon$, $h$ in Eq. (3). The first is responsible for the average parity of elementary cycles, and the second is responsible for the relative prevalence of friendly and unfriendly relationships. The normalization factors $1/n_{t}$ and $1/n_{e}$ in front of the sums (3) are chosen so that both terms $P$ and $M$ are extensive in $N$. The partition function (2) describes a thermodynamic system with temperature $T=1/\beta$. The temperature sets the scale for the coupling constants $\varepsilon$ and $h$. The partition function
@@
depends on $\varepsilon,h$ and $T$ via effective parameters $\varepsilon^{\prime}=\varepsilon/T$, $h^{\prime}=h/T$. In the limit $\varepsilon^{\prime}\rightarrow\infty$, the system is structurally balanced. When the sum $\{s_{ab}=\pm 1\}$ in Eq. (4) is restricted to the subspace of balanced states $\{s_{ab}=\pm 1\}_{B}$, such that $s_{ab}s_{bc}s_{ca}=1$ for all triangles, the partition function reduces to
@@
which can be obtained by the Hubbard–Stratonovich transformation applied to $\exp\left(h^{\prime}/n_{e}(\sum_{a}\sigma_{a})^{2}\right)$ in Eq. (7). The integral representation is particularly useful in the thermodynamic limit $N\rightarrow\infty$ to determine the asymptotic behavior of the partition function, which, in the leading order, grows exponentially with the coefficient
@@
For $h^{\prime}\leq 1/2$, the last equation gives $x_{*}=0$, while for $h^{\prime}>1/2$, it yields positive $x_{*}$, which for small positive $\Delta h=h^{\prime}-1/2$ behaves as $x_{*}=\sqrt{24}\sqrt{\Delta h}+o(\sqrt{\Delta h})$. The second derivative $\partial^{2}_{h^{\prime}h^{\prime}}\phi_{I}$ is discontinuous for $h^{\prime}=1/2$, where it jumps from 0 to $6$. The result is shown in Fig. 1. The limit curve for $N\rightarrow\infty$ is obtained using Eqs. (11) and (12) while the finite size results are obtained from Eq. (8).
@@
A comparison of the Monte Carlo results with the theoretical predictions resulting from the equivalence of the Heider model for $\varepsilon^{\prime}\rightarrow\infty$ with the Ising model is presented in Fig. 1. Both models give exactly the same results. One can also see that the normalized variance $\chi_{m}$ is a function of the field $h^{\prime}$, which is a good scaling variable in the limit $\varepsilon^{\prime}\rightarrow\infty$. On the other hand, for $\varepsilon^{\prime}=0$, the normalized variance can easily be found to be $\chi_{m}=\cosh^{-2}(h^{\prime}/n_{e})/n_{e}^{2}$. This follows from the independence of edge signs for $\varepsilon^{\prime}=0$. So we see that in this case $h^{\prime}/n_{e}$ rather than $h^{\prime}$ is a good scaling variable. The corresponding curve $\chi_{m}$ plotted as a function of $h^{\prime}$ will broaden as $N$ increases.
The situation is illustrated in Fig. 2 where we plot $\chi_{m}$ vs. $h^{\prime}$ for different values of $\varepsilon^{\prime}$ for $N=20$. The curves for $\varepsilon^{\prime}\geq 6$ have comparable widths and a more or less constant position. They only slightly deviate from the limiting curve, while the curves for $\varepsilon^{\prime}\leq 3$ are much wider and have a moving mode position. A similar picture is observed for other sizes of the system $N$.
Figure 2: Evolution of the normalized variance $\chi_{m}$ with $\varepsilon^{\prime}$, for $N=20$. The curves are obtained by Monte Carlo simulations. For $N=20$, the critical value is $\varepsilon^{\prime}_{cr}\approx 6$, see Fig. 4. For $\varepsilon^{\prime}>\varepsilon^{\prime}_{cr}$, the curves quickly converge to a limit curve. For $\varepsilon^{\prime}<\varepsilon^{\prime}_{cr}$, the curves broaden when $\varepsilon^{\prime}$ increases; they also broaden when NN increases although this is not shown in the figure
The picture is correlated with the phase diagram of the model. The model undergoes a first order phase transition at a critical value of $\varepsilon^{\prime}$, where the average parity of elementary loops $p$ has a discontinuity when plotted as a function of $\varepsilon^{\prime}$. In the case of finite $N$, instead of a discontinuity, a sharp increase is observed with a slope of $1/N$ in a narrow interval. At the transition point, the histogram of $p$ has a bimodal structure consisting of two peaks corresponding to coexisting phases. The appearance of a bimodal histogram is a strong indication of a first-order phase transition. When the value $\varepsilon^{\prime}$ changes, so do the proportions of the peaks. There are different ways to define the pseudo-critical value [18, p. 162]. We find it convenient to define it as the value of $\varepsilon^{\prime}$ at which the mean of the distribution of $p$ is equal to $1/2$, which only occurs when the two phases coexist (in comparable proportions).
@@
Using this definition, we found numerically that the pseudo-critical value grows linearly with $N$ roughly as: $\varepsilon^{\prime}_{cr}\approx(N-2)/3$; see Fig. 4. The estimated dependence on $N$ does not differ significantly from the mean field result, which also predicts that for large $N$ the pseudo-critical value increases linearly with $N$ as $\varepsilon^{\prime}\approx 0.286(N-1)$ for $h^{\prime}=0$ [19].
Figure 4: The pseudo-critical value $\varepsilon^{\prime}_{cr}$ for different values of $N$, for $h^{\prime}=0$. The phase transition is first order. At the phase transition two phases co-exist: one with $p<1/2$ and the other with $p>1/2$, see Fig. 3. The pseudo-critical values $\varepsilon^{\prime}_{cr}$ shown as points for different NN, are obtained from the condition that the mean (the center of mass of the histogram) is $\langle p\rangle=1/2$. The line $\varepsilon^{\prime}_{cr}=(N-2)/3$, drawn as a solid line, passes through the points
@@
The equivalence of the balanced Heider states with the Ising states is a direct consequence of Eq. (6), which resolves simultaneously the constraints $\pi(\gamma_{e})=1$ for all elementary cycles. This equivalence holds for the Heider model on any graph, not only on the complete graph. The phase transition in the Heider model, for balanced states, is inherited from the phase transition of the Ising model. The order of this transition thus depends on the dimensionality of the problem. In particular, the susceptibility $\chi_{m}=\partial^{2}_{h^{\prime}h^{\prime}}\ln Z$ exhibits a logarithmic divergence at the critical value of the field in two dimensions [16], and a power-law divergence $\chi_{m}\sim|\Delta h|^{-\alpha}$, with $\alpha=0.11008708(35)$, inherited from the three dimensional Ising model [20].

### v2

@@ I Introduction
The problem of Heider balance is often discussed in connection with homophily, i.e., the tendency of a node to form a positive link with a node of similar characteristics [6, 7, 8]. The outcome of this combination can be complex. Reference 9 demonstrates the competition between the homophily-driven formation of groups and the local adjustment towards balance. In Reference 10, the system ability to achieve the balance depends on the number $G$ of attributes that determine the homophily process. It is demonstrated that above some critical value of $G$ depending on the size of the system $N$, the phase termed ‘paradise’ (all positive links) is attained. A similar competition between homophily and balance has been considered in Reference 11, with somewhat different results on the $G$-dependence of the final state; for large $G$, the balance is destroyed. We also note that an increase in social connectivity has been found to produce fragmentation and polarization [12, 13]. This discussion, apparently relevant for mathematical social sciences, seems far from a final conclusion.
Signed networks belong to a large family of complex systems, where a prescribed time evolution produces some stationary states. In undirected networks, the evolution is directed as to minimize a work function, analogous to energy in physical systems. A physical system close to the network considered here is a spin glass [14, 15, 16]. In some models, signs of the exchange integrals of the magnetic interaction between nodes are randomly distributed. The key difference between our signed networks and a spin glass is that in the latter, magnetic moments $\pm 1$ of atoms at the network nodes undergo the evolution, while the links are fixed. On the contrary, in the signed network discussed here the dynamics applies to links, and spins are absent. Despite decades of efforts, physics of spin glasses remains a fertile field of research [17]. From this perspective, signed networks promise a path to understanding ‘a simpler complex system’.
On the other hand, sociological attempts to simulate complex networks are spread on a broader front of research [18]. Still a tendency can be observed to focus on relations between individuals [19, 20], rather than on individuals or groups as social entities. In computational models, the key effort is to identify a social process, and—if possible—a work function which could be minimized. This program can be executed only separately for a given social phenomenon. The present text is devoted to minimization of cognitive dissonance that drives the system to a balanced state.
@@
The analysis of the equivalence of the Heider model and the Ising model [28, 29, 30, 31] emphasizes the importance of spin models in the description of social systems. The Ising model is simple enough to be grasped and discussed in detail, both analytically and numerically. At the same time, it generates a reach spectrum of phenomena allowing to better understand: models of opinion dynamics [32, 33, 34, 35]; social segregation [36, 37]; crowd dynamics [38, 39]; language dynamics [40, 41]; games theory [42, 43]; tax evasion [44, 45]; modeling elections [46, 47]; large-scale social behaviors [48, 49]; homophily in social relations [10, 50]; psychological theories evaluation [51]; immigration, integration and ghetto formation [52]; the spread of infectious diseases [53, 54] (see References 55, 56 for comprehensive reviews), to mention only several in the field of sociophysics [57, 58, 59, 60, 61].
The concept of the applied field is useful in social psychology [62, 63], as it can represent incentives to improve or deteriorate interpersonal relationships. An external field in this role has been integrated into the Heider model in References 64, 65. The field and thermal noise complement the toolbox of statistical mechanics [66]. Our aim here is to demonstrate the equivalence of the balanced system in the presence of an external field and the standard Ising system without a field.
II Methods
@@
A balanced network has the following topological property [2]: the parity $\pi$ of any path $\gamma=(a_{1}a_{2},a_{2}a_{3},\ldots,a_{n}a_{n+1})$, defined as the product of the signs of the edges that form the path
$\pi(\gamma)=\prod_{i=1}^{n}a_{i}a_{i+1}$ (1)
@@
where $n_{t}=N_{t}/N=(N-1)(N-2)/6$ is the density of triangles per node, and $n_{e}=N_{e}/N=(N-1)/2$ is the density of edges per node. The first sum runs over all triangles, and the second one runs over the links of the network. In physical terms, $M$ can be called the edge magnetization, and $m=M/N$ the magnetization density. Positive/negative $M$ indicates friendly/unfriendly relations. The term $P$ sums the parities of the elementary cycles; $p=P/N$ is the average parity per cycle. There are two coupling constants $\varepsilon$, $h$ in Equation (3). The first is responsible for the average parity of elementary cycles, and the second is responsible for the relative prevalence of friendly and unfriendly relationships. The normalization factors $1/n_{t}$ and $1/n_{e}$ in front of the sums (3) are chosen so that both terms $P$ and $M$ are extensive in $N$. The partition function (2) describes a thermodynamic system with temperature $T=1/\beta$. The temperature sets the scale for the coupling constants $\varepsilon$ and $h$. The partition function
@@
depends on $\varepsilon,h$ and $T$ via effective parameters $\varepsilon^{\prime}=\varepsilon/T$, $h^{\prime}=h/T$. In the limit $\varepsilon^{\prime}\rightarrow\infty$, the system is structurally balanced. When the sum $\{s_{ab}=\pm 1\}$ in Equation (4) is restricted to the subspace of balanced states $\{s_{ab}=\pm 1\}_{B}$, such that $s_{ab}s_{bc}s_{ca}=1$ for all triangles, the partition function reduces to
@@
which can be obtained by the Hubbard–Stratonovich transformation applied to $\exp\left(h^{\prime}/n_{e}(\sum_{a}\sigma_{a})^{2}\right)$ in Equation (7). The integral representation is particularly useful in the thermodynamic limit $N\rightarrow\infty$ to determine the asymptotic behavior of the partition function, which, in the leading order, grows exponentially with the coefficient
@@
For $h^{\prime}\leq 1/2$, the last equation gives $x_{*}=0$, while for $h^{\prime}>1/2$, it yields positive $x_{*}$, which for small positive $\Delta h=h^{\prime}-1/2$ behaves as $x_{*}=\sqrt{24}\sqrt{\Delta h}+o(\sqrt{\Delta h})$. The second derivative $\partial^{2}_{h^{\prime}h^{\prime}}\phi_{I}$ is discontinuous for $h^{\prime}=1/2$, where it jumps from 0 to $6$. The result is shown in Figure 1. The limit curve for $N\rightarrow\infty$ is obtained using Equations (11) and (12) while the finite size results are obtained from Equation (8).
@@ III Results
IV Discussion
A comparison of the Monte Carlo results with the theoretical predictions resulting from the equivalence of the Heider model for $\varepsilon^{\prime}\rightarrow\infty$ with the Ising model is presented in Figure 1. Both models give exactly the same results. One can also see that the normalized variance $\chi_{m}$ is a function of the field $h^{\prime}$, which is a good scaling variable in the limit $\varepsilon^{\prime}\rightarrow\infty$. On the other hand, for $\varepsilon^{\prime}=0$, the normalized variance can easily be found to be $\chi_{m}=\cosh^{-2}(h^{\prime}/n_{e})/n_{e}^{2}$. This follows from the independence of edge signs for $\varepsilon^{\prime}=0$. So we see that in this case $h^{\prime}/n_{e}$ rather than $h^{\prime}$ is a good scaling variable. The corresponding curve $\chi_{m}$ plotted as a function of $h^{\prime}$ will broaden as $N$ increases.
The situation is illustrated in Figure 2 where we plot $\chi_{m}$ vs. $h^{\prime}$ for different values of $\varepsilon^{\prime}$ for $N=20$. The curves for $\varepsilon^{\prime}\geq 6$ have comparable widths and a more or less constant position. They only slightly deviate from the limiting curve, while the curves for $\varepsilon^{\prime}\leq 3$ are much wider and have a moving mode position. A similar picture is observed for other sizes of the system $N$.
Figure 2: Evolution of the normalized variance $\chi_{m}$ with $\varepsilon^{\prime}$, for $N=20$. The curves are obtained by Monte Carlo simulations. For $N=20$, the critical value is $\varepsilon^{\prime}_{cr}\approx 6$, see Figure 4. For $\varepsilon^{\prime}>\varepsilon^{\prime}_{cr}$, the curves quickly converge to a limit curve. For $\varepsilon^{\prime}<\varepsilon^{\prime}_{cr}$, the curves broaden when $\varepsilon^{\prime}$ increases; they also broaden when NN increases although this is not shown in the figure
The picture is correlated with the phase diagram of the model. The model undergoes a first order phase transition at a critical value of $\varepsilon^{\prime}$, where the average parity of elementary loops $p$ has a discontinuity when plotted as a function of $\varepsilon^{\prime}$. In the case of finite $N$, instead of a discontinuity, a sharp increase is observed with a slope of $1/N$ in a narrow interval. At the transition point, the histogram of $p$ has a bimodal structure consisting of two peaks corresponding to coexisting phases. The appearance of a bimodal histogram, presented in Figure 3, is a strong indication of a first-order phase transition. When the value $\varepsilon^{\prime}$ changes, so do the proportions of the peaks. There are different ways to define the pseudo-critical value [67, p. 162]. We find it convenient to define it as the value of $\varepsilon^{\prime}$ at which the mean of the distribution of $p$ is equal to $1/2$, which only occurs when the two phases coexist (in comparable proportions).
@@ IV Discussion
Using this definition, we found numerically that the pseudo-critical value grows linearly with $N$ roughly as: $\varepsilon^{\prime}_{cr}\approx(N-2)/3$; see Figure 4. The estimated dependence on $N$ does not differ significantly from the mean field result, which also predicts that for large $N$ the pseudo-critical value increases linearly with $N$ as $\varepsilon^{\prime}\approx 0.286(N-1)$ for $h^{\prime}=0$ [68].
Figure 4: The pseudo-critical value $\varepsilon^{\prime}_{cr}$ for different values of $N$, for $h^{\prime}=0$. The phase transition is first order. At the phase transition two phases co-exist: one with $p<1/2$ and the other with $p>1/2$, see Figure 3. The pseudo-critical values $\varepsilon^{\prime}_{cr}$ shown as points for different NN, are obtained from the condition that the mean (the center of mass of the histogram) is $\langle p\rangle=1/2$. The line $\varepsilon^{\prime}_{cr}=(N-2)/3$, drawn as a solid line, passes through the points
@@
Although equivalence between Heider and Ising models may seem trivial, its consequences are not:
  • First of all, for $h=0$, the model has local symmetry. This type of symmetry is highly non-trivial and in the language of field theory it is called gauge symmetry. Using this equivalence, one can compute the volume of the symmetry group (gauge orbit).
  • Second, for $h\neq 0$, local symmetry is broken, but the equivalence allows us to show that the partition function of the Heider model can be mapped onto the partition function of the Ising model, in which the social field $h$ corresponds to the reciprocal of the temperature, and the susceptibility in the Heider model corresponds to the specific heat in the Ising model due to the social field. To the best of our knowledge, the equivalence of these two models was not known.
The equivalence of the balanced Heider states with the Ising states is a direct consequence of Equation (6), which resolves simultaneously the constraints $\pi(\gamma_{e})=1$ for all elementary cycles. This equivalence holds for the Heider model on any graph, not only on the complete graph. The phase transition in the Heider model, for balanced states, is inherited from the phase transition of the Ising model. The order of this transition thus depends on the dimensionality of the problem. In particular, the susceptibility $\chi_{m}=\partial^{2}_{h^{\prime}h^{\prime}}\ln Z$ exhibits a logarithmic divergence at the critical value of the field in two dimensions [65], and a power-law divergence $\chi_{m}\sim|\Delta h|^{-\alpha}$, with $\alpha=0.11008708(35)$, inherited from the three dimensional Ising model [69].